if the perimeter of the rectanglangle is 48 yards find the length if it is 3x the width
Answer by checkley75(3666)

2X+2*3X=48
2X+6X=48
8X=48
X=48/8
X=6 ANSWER FOR THE LENGTH.
3*6=18 ANSWER FOR THE WIDTH.
PROOF
2*6+2(3*6)=48
12+2*18=48
12+36=48
48=48